# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ V Discussion
Our results demonstrate that the Laplacian and RBF kernels can be used to obtain accurate and efficient estimates of OTOCs and related quantities, such as the lower bound in (17). The main cost involved in this approach is that of producing the training dataset. Once we have a trained model, however, it can be used to make predictions for OTOC values instead of continued implementations of the MPO-based algorithm. One could also imagine that if one had access to a fault-tolerant QC, it could be used to supply training data for a ML model. Using KMs, as in this work, would then allow one to obtain an approximation of the correlation landscape in some subset of the Hamiltonian parameter space, without requiring further (potentially expensive) use of the QC. In either case, having the ability to quickly generate approximate predictions of OTOC values for a wide range of Hamiltonians could, for example, assist in identifying interesting behaviours for particular sets of Hamiltonian parameters, which one could then validate by other means.
Given the cost of producing training data, it is natural to ask whether one can obtain similar results with fewer data. To investigate this, we perform a similar procedure as above, but with a varying amount of training data. First, we fix the hyperparameter values for each kernel in each problem instance to be those which were chosen during the cross-validation performed on all of the training data. Next, we choose different training dataset sizes $M\in\mathbb{N}$ to supply to the models. Specifically, for each $M\in\{50m:m=1,2,\ldots,20\}$, we randomly sample $M$ of the 1000 training data points, and repeat this 20 times to obtain 20 (possibly intersecting) subsets of size $M$. We then train a model on each of these 20 subsets of the training data. Finally, we calculate the average $R^{2}$ score for the 20 models making predictions on the entire testing set, as an indicator of how the models perform with only $M$ training data samples. The results of this analysis are shown in Figure 14 for the 40-qubit problem instances. The plots show that, even with only a few hundred training data points, the average $R^{2}$ scores for the Laplacian and RBF kernels often remain above $0.8$, and occasionally reach above $0.9$. Thus, the kernels are still capable of learning from fewer training data samples than were supplied in the earlier analysis.
Figure 14: Average coefficient of determination $R^{2}$ for each kernel making predictions on the 40-qubit testing sets, plotted against the number of training data samples $M$ used to train the associated ML models. Each point in the plots corresponds to the numerical value of $R^{2}$ averaged over 20 different models using the same kernel. Each of the models is the result of applying Algorithm 1 with a random subset of the training data containing $M$ of the total 1000 training data samples. The hyperparameters are fixed to be the best hyperparameters found during the 10-fold cross-validation performed on all 1000 training data samples. Plots (a), (b), (c) and (d) show the average $R^{2}$ score on the testing datasets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Plots (e), (f), (g) and (h) show the average $R^{2}$ score on the testing datasets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively.
Another point we wish to highlight is that the KM approach allows us to obtain a good approximation of the various OTOC functions (as calculated by the MPO-based algorithm) for systems evolving up to a time linear in the system size. In [54], the authors applied restricted Boltzmann machines to learn OTOCs at early times, which they refer to as the values of $t$ for which $\textrm{Re}[F(t)]>0.85$. In contrast, one can infer from Figure 4 that the OTOCs occurring in $\mathscr{O}_{Sum}$ have decayed far below 0.85 (as does the XZ-OTOC). This shows that the KM approach is capable of predicting OTOCs for larger timescales than was achieved in [54].
@@
One limitation of the approach taken here is that the OTOC labels, which were generated using the MPO-based algorithm, are only approximate due to the finite bond dimension. As can be seen in Figure 5, the maximum bond dimension $\chi$ selected for producing some of the data would ideally have been larger, especially for the data produced with $\mathscr{H}_{4}$ (see Figures 5(d) and 5(h)). However, we did not have the computational resources to increase $\chi$ further in these cases. On the other hand, for $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ and $\mathscr{H}_{3}$, the values of $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ appear to not be changing significantly upon further increase of the bond dimensions above those used to produce the datasets. We also note that the plots in Figure 5 correspond to the 40-qubit systems, which is the largest of the systems used here. For the smaller systems we considered, the convergence with respect to $\chi$ will occur more quickly.
Another potential limitation of our ML approach is that one should not expect the kernels to be able to extrapolate beyond the ball in $\mathbb{R}^{3}$ from which the parameter values were sampled. Recall, the testing data, on which we evaluated the performance of the models, was drawn from the same distribution as the training data. Thus, what we have demonstrated is that the kernels are able to make accurate predictions for other Hamiltonians within this region in parameter space. This could nevertheless be useful since this region may contain Hamiltonians which are theoretically interesting but difficult to prepare physically. Indeed, as seen in Figure 4, these regions for different sets of Hamiltonians are sufficiently large to capture a qualitatively rich variety of behaviours in the underlying quantum systems.
@@ VI Conclusion
In this paper, we explored the application of classical KMs to learn and predict the OTOC, an important quantity used to investigate quantum information scrambling. By viewing the XZ-OTOC, and the sum of OTOCs in (17), as functions of parameters associated with a parameterised set of Hamiltonians, we frame the learning task as a regression problem and employ 6 standard kernel functions. For all 4 of the parameterised sets of Hamiltonians considered, the results show that the Laplacian and RBF kernels perform best. They are both capable of closely approximating the XZ-OTOC and the sum of OTOCs (as calculated by the MPO-based algorithm) from a moderate amount of training data. Even for systems with reasonably large qubit numbers, the Laplacian and RBF kernels often achieve $R^{2}$ scores on the testing data between $0.9$ and $0.99$, and almost always at least 0.8, accompanied by small values for the RMSE and MAE.
With these trained kernels, it is possible to efficiently investigate the behaviour of the OTOC across a wide range of systems. Of course, obtaining the trained kernels required significant computational costs to produce training data. However, we have shown here that it is possible to obtain a sufficient amount of data to accurately learn the OTOC functions that we considered. Indeed, we saw that KMs can achieve good learning performance metrics (e.g., $R^{2}>0.8$) from a few hundred data points. The value of the KM approach, therefore, is in situations where one would like to know the behaviour of the OTOCs across a much larger number of systems ($\gtrsim 10^{3}$) within the parameterised sets. The KMs can efficiently predict the values for all of these systems from a few hundred data points, which is a comparatively much smaller cost than simulating each of these systems individually.
Our work improves on aspects of previous work [54] which used ML to approximate OTOCs for early times. Here we derive ML models which can make accurate predictions over timescales which are linear in the size of the underlying quantum system, exceeding the early timescales considered in [54]. We also provide a significant body of numerical evidence showing that the KM approach works well for many instances of the problem. Nonetheless, one should not expect that the KM approach will allow us to predict OTOCs and related functions for arbitrary inputs, but only for those sampled from the same subset as the input training data. However, this can still be useful since the domains of the learned models are sufficiently large to describe a diverse spectrum of behaviours in a quantum system.
@@
The paper naturally lends itself to a handful of possible extensions, which we suggest as future research directions. For example, one may like to investigate whether the KM approach works with other OTOCs, other linear combinations of OTOCs, or other types of correlation functions such as the Loschmidt echo [78]. We could also examine how the approach works for different or larger parameterised sets of Hamiltonians. For instance, we could investigate sets with $d>3$ (i.e., parameterised by 4 or more parameters), sets containing Hamiltonians describing quantum systems in higher dimensions, or sets containing highly non-local Pauli strings in one or more terms. Though producing data for the latter two cases would likely be challenging. Similarly, it might be interesting to apply quantum kernel methods [79, 80, 81] to investigate whether a problem-specific quantum kernel may offer significant increases in learning performance. Finally, it could be useful to explore how the results change in response to larger values of $\chi$ in the MPO calculations, and increases in the size of the region from which we sample the input data.
@@
The datasets produced in this study are available in the GitHub repository at https://github.com/John-J-Tanner/ITensor-OTOC-Data. The repository includes all relevant datasets necessary to replicate the study’s results. For further information or inquiries, please contact the corresponding author, John Tanner, at john.tanner@uwa.edu.au.
@@ Appendix A Supplementary results
Appendix B Hyperparameter values
@@
Table I: Lists of hyperparameters values which were trialled during the 10-fold cross-validation performed on the training datasets. The hyperparameter $\lambda$ refers to the regularisation strength which appears in (1), while the hyperparameters $\gamma$, $c_{0}$ and $d$ correspond to those which appear in definitions of the kernels in (24), (25), (26), (27), (28) and (29).
@@
Table II: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{1}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all 10 folds of the training dataset.
@@
Table III: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{2}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all 10 folds of the training dataset.
@@
Table IV: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{3}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all 10 folds of the training dataset.
@@
Table V: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{4}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all 10 folds of the training dataset.
Appendix C Numerical results
C.1 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{1}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 0.7886 0.8122 0.8156 0.8214 0.8156 0.8147 0.8030 0.7941 MAE 0.7091 0.7296 0.7275 0.7307 0.7186 0.7159 0.6991 0.6857 $R^{2}$1.9284 -4.1339 -3.5942 -3.7616 -3.4297 -3.3197 -3.0432 -2.8091 Polynomial RMSE 0.1722 0.1493 0.1755 0.1563 0.1669 0.1694 0.1716 0.1775 MAE 0.1228 0.1076 0.1185 0.1121 0.1200 0.1191 0.1249 0.1274 $R^{2}$ 0.8604 0.8265 0.7872 0.8276 0.8145 0.8132 0.8154 0.8097 RBF RMSE 0.0867 0.0837 0.1256 0.1212 0.1359 0.1216 0.1191 0.1117 MAE 0.0431 0.0489 0.0586 0.0772 0.0790 0.0739 0.0674 0.0594 $R^{2}$ 0.9646 0.9454 0.8910 0.8963 0.8771 0.9038 0.9111 0.9246 Laplacian RMSE 0.1137 0.0625 0.1057 0.0801 0.0983 0.0990 0.0978 0.0871 MAE 0.0763 0.0388 0.0495 0.0463 0.0529 0.0660 0.0636 0.0559 $R^{2}$ 0.9391 0.9696 0.9228 0.9547 0.9356 0.9362 0.9401 0.9542 Sigmoid RMSE 0.2616 0.1322 0.2259 0.1601 0.1669 0.1538 0.2240 0.2441 MAE 0.1852 0.0895 0.1617 0.1123 0.1176 0.1081 0.1696 0.1899 $R^{2}$ 0.6777 0.8639 0.6475 0.8191 0.8144 0.8461 0.6852 0.6400 Cosine RMSE 0.7916 0.8117 0.8135 0.8218 0.8153 0.8139 0.8066 0.7971 MAE 0.7116 0.7285 0.7240 0.7306 0.7175 0.7136 0.7024 0.6882 $R^{2}$1.9508 -4.1287 -3.5710 -3.7659 -3.4271 -3.3113 -3.0796 -2.8382 Training Linear RMSE 0.7692 0.7919 0.7971 0.8052 0.8025 0.7979 0.7870 0.7745 MAE 0.6823 0.6985 0.6996 0.7035 0.6971 0.6893 0.6750 0.6589 $R^{2}$1.7147 -3.4479 -3.1980 -3.1767 -2.9977 -2.8207 -2.6211 -2.4198 Polynomial RMSE 0.1519 0.1329 0.1398 0.1414 0.1439 0.1426 0.1520 0.1583 MAE 0.1071 0.0939 0.0992 0.1006 0.1035 0.0997 0.1078 0.1110 $R^{2}$ 0.8942 0.8747 0.8709 0.8712 0.8715 0.8779 0.8650 0.8572 RBF RMSE 0.0077 0.0334 0.0138 0.0809 0.0667 0.0723 0.0580 0.0476 MAE 0.0049 0.0224 0.0078 0.0557 0.0452 0.0491 0.0390 0.0308 $R^{2}$ 0.9997 0.9921 0.9987 0.9578 0.9724 0.9686 0.9803 0.9871 Laplacian RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0075 0.0059 0.0047 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0054 0.0042 0.0033 $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9998 0.9999 Sigmoid RMSE 0.2539 0.0912 0.2056 0.1380 0.1423 0.1204 0.2177 0.2544 MAE 0.1791 0.0666 0.1470 0.0996 0.1009 0.0881 0.1567 0.1892 $R^{2}$ 0.7042 0.9410 0.7206 0.8773 0.8743 0.9130 0.7229 0.6310 Cosine RMSE 0.7690 0.7923 0.7979 0.8052 0.8027 0.7983 0.7860 0.7736 MAE 0.6816 0.6983 0.6983 0.7030 0.6961 0.6873 0.6746 0.6578 $R^{2}$1.7129 -3.4530 -3.2072 -3.1761 -2.$9$2.8250 -2.6125 -2.4122
@@
Table VI: Learning performance metrics for the models trained with the hyperparameters listed in Table II for $\mathscr{O}_{XZ}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{1}$ with labels determined by $\mathscr{O}_{XZ}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 7, though not all of them are visible within the chosen plot range.
C.2 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{1}$

### v2

@@ V Discussion
The results in the previous section demonstrate that the Laplacian and RBF kernels can be used to obtain accurate estimates of OTOCs and related quantities, such as the lower bound in (18), from a relatively small amount of training data. The suitability of these kernels for the learning problems described in Section III.2 suggests that the OTOC functions $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ change smoothly in such a way that the value of either function, for a given input, can be closely inferred from the function values at nearby inputs. This behaviour is captured well by the Laplacian and RBF kernels since the value of these kernels decay exponentially with the distance between their arguments (as measured by the $l_{1}$ and squared $l_{2}$ norms respectively). The polynomial and sigmoid kernels also performed reasonably well, but the results show that the Laplacian and RBF kernels consistently performed best of all six kernels that we trialled in this work.
Given that the Laplacian and RBF kernels are capable of accurately predicting OTOCs from a small amount of training data, this means that we can use KMs together with the MPO-based algorithm [42] to obtain a reduction in overall runtime when one is interested in evaluating the OTOC for many different systems and values of time. Specifically, once we have a trained model, it can be used to substitute the MPO-based algorithm by making accurate predictions of OTOC values in a time scaling linearly in just the size of the training dataset (i.e., independent of the system size).
To see this, suppose that we are interested in evaluating $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$ for a large number $N\in\mathbb{N}$ of different inputs drawn from the spherical domains described in Section III.1. From (32), the runtime scaling of using the MPO-based algorithm to calculate these values is
$\mathscr{O}\left(Nn\chi^{3}t/\Delta t+Nn\chi^{4}\right).$ (35)
If, however, we decide to instead produce $M$ training data samples using the MPO-based algorithm, where $M<<N$, and then predict the OTOC values using the Laplacian or RBF kernels, then the overall runtime scales as
$\mathcal{O}\left(Mn\chi^{3}t/\Delta t+Mn\chi^{4}+N_{\lambda}N_{\gamma}M^{3}+NM\right),$ (36)
where $N_{\lambda}$ and $N_{\gamma}$ are the number of different values of $\lambda$ and $\gamma$, respectively, used in the grid search. The first two terms in (36) come from producing the $M$ training datapoints, the third term from training the model using KMs with a cross-validation by grid search, and the last term from predicting the OTOC values with the trained model (see the end of Section II.2). Comparing (35) with (36) shows that, after producing the training dataset and training the model, the cost of making further predictions is reduced significantly from $\mathcal{O}\left(n\chi^{3}t/\Delta t+n\chi^{4}\right)$ to $\mathcal{O}(M)$ per prediction. So in cases where the number of required training data samples is much smaller than the overall number of predictions we would like to make (i.e., $M<<N$), applying KMs may significantly reduce the overall computational cost of obtaining the OTOC values.
This may be useful, for example, to a user who is interested in finding a quantum system which exhibits some interesting property that can be inferred from OTOCs. For instance, perhaps one is interested in seeking a highly chaotic system, so that local information quickly spreads across the system, which may be convenient in situations such as the classical communication protocol depicted in Figure 2. One possible way of finding such a system would be to search for a choice of Hamiltonian parameters which results in the greatest quantum Lyapunov exponent. This entails calculating the short-time regime behaviour of the OTOC for many different systems and values of time associated with various choices of the Hamiltonian parameters. In such cases, where $N>>M$, the proposed runtime reduction could be practically useful.
The main cost of applying KMs to the learning problems we consider is that of producing the training datasets. And in order to produce the data in this work, as discussed in Section III.4, we make use of the MPO-based algorithm, since this algorithm lends itself well to simulating local 1D quantum systems. However, producing the data, even with this algorithm, can become expensive as the size of the system becomes large. This is because the maximum evolution time that we simulate, and often the necessary maximum bond dimension, increases with system size, leading to a high degree polynomial runtime scaling in the system size.
One could also consider generating data, or even directly calculating the OTOC values, on a QC instead. However, we expect that currently available NISQ devices will require substantial improvements before it will be possible to match the capabilities of tensor network methods for simulating 1D quantum systems. For example, neutral atom processors provide large qubit numbers, but exhibit two qubit gate error rates of roughly 0.5% [3], which severely limits circuit depths. In contrast, trapped ion devices provide small two qubit gate errors around 0.1%, but only have on the order of 50 qubits [80]. In fact, of all currently available devices, only IBM’s superconducting processors [81] have enough qubits with sufficiently small error rates to challenge classical methods for 1D systems. And it appears that even these devices would not be capable of generating the datasets in this work with good accuracy.
To see this, in Section I.B.1 of [82], the authors state that performing quantum simulations on NISQ hardware typically requires circuit depths greater than what can be performed with current devices, even those achieving a 0.1% average two qubit gate error rate. We believe the quantum simulations required to calculate the OTOCs considered in our work are no exception. For example, Trotterising the Hamiltonians discussed in Section III.1 using a Trotter step size of $\Delta t=0.05$, up to a time equal to the system size $n$, requires at least $20n$ layers of two qubit gates. The number of layers which can feasibly be implemented depends on many factors. But as a baseline, in 2023 the authors of [81] reported an implementation of 60 layers of two qubit CNOT gates with high fidelity on IBM’s 127-qubit superconducting processor ibm_kyiv. Since the publication of [81], the baseline error levels have marginally improved, so it is likely that more than 60 layers of CNOT gates would now be possible. However, the current error rates still exceed those which are expected to be necessary for many useful near-term applications (see Table I of [82]). And when we compare the baseline of 60 layers with $20n$, it becomes apparent that the number of layers required to use $\Delta t=0.05$ will quickly exceed that which is feasible.
One could argue that we may simply increase the Trotter step size to reduce the number of layers, while still simulating up to a time $t_{max}=n$. However, this will result in the Trotter error $\mathcal{O}\left((\Delta t)^{3}\right)$ scaling cubically with the system size. For example, with 40 qubits and 100 layers, we would have $\Delta t=0.4$, which increases the magnitude of the Trotter error by a factor of $\sim 500$ when compared with $\Delta t=0.05$. This leads us to believe that current NISQ devices are not yet capable of generating the data used in this work with the same accuracy as the MPO-based algorithm, especially in cases with large $n$. This is not in contradiction with the claims of [81, 80], since there the authors simulate quantum systems in higher dimensions, while here we consider only 1D quantum systems, to which tensor network methods are well suited. Instead, we simply argue that classical tensor networks are preferable over methods utilising NISQ devices for generating the datasets used in this work.
@@
Despite this, one of the limitations of the approach taken here is that the OTOC labels, which were generated using the MPO-based algorithm, are only approximate due to the use of a finite bond dimension. The analysis presented in Figure B.1 suggests that the chosen maximum bond dimension $\chi$ was sufficiently large for most problem instances, since the OTOC values appear to have converged. However, calculations for some of the Hamiltonians from the set $\mathscr{H}_{4}$ (see Figures B.1(d) and B.1(h)) could have benefitted from a larger $\chi$, though this would have been beyond our computational resources. We also note that the plots in Figure B.1 correspond to the 40-qubit systems, which is the largest of the systems used here. For smaller system sizes, the convergence with respect to $\chi$ will occur more quickly.
Another limitation of our ML approach that warrants discussion is that one should not expect the kernels to be able to extrapolate beyond the ball in $\mathbb{R}^{3}$ from which the input training data is sampled. Recall, the testing data, on which we evaluated the performance of the models, was drawn from the same distribution as the training data. Thus, what we have demonstrated is that the kernels are able to make accurate predictions for other Hamiltonians within this region in parameter space. This may still be useful however, since the region of parameter space on which the models perform well captures evolution times which scale linearly in the system size. As discussed in Section II.3, we expect that this captures the short-time regime, and possibly part of the intermediate-time regime. Accordingly, we expect that the models may be used to make predictions for sufficiently large values of times to be of theoretical interest. In contrast, the only related work of which we are aware [55] uses ML techniques to learn OTOCs in just a small part of the short-time regime (when Re$[F_{jk}](t)<0.85$). However, a direct comparison of our work with [55] is not straightforward since they demonstrated their method using a 2D quantum system, while this work focuses on 1D systems.
@@ VI Conclusion
In this paper, we explored the application of classical KMs to learn and predict the OTOC, an important quantity used to investigate quantum information scrambling. By viewing the XZ-OTOC, and the sum of OTOCs in (18), as functions of parameters associated with a parameterised set of Hamiltonians, we frame the learning task as a regression problem and employ six standard kernel functions. For all four of the parameterised sets of Hamiltonians considered, each describing local 1D quantum systems, we provide a significant body of numerical results (see Appendix D) which shows that the Laplacian and RBF kernels perform best. Indeed, both kernels closely approximate the XZ-OTOC and the sum of OTOCs, achieving $R^{2}$ scores on the testing data of at least 0.7167 and typically around 0.9, with small RMSE and MAE (see Table I), from a small number of training datapoints and a standard cross-validation. In fact, Figure 5 shows that the kernels can achieve $R^{2}$ scores in excess of 0.8 even with just a few hundred datapoints.
Since the kernels perform well with a small amount of training data, we can apply KMs to reduce the overall computational cost of extensively evaluating OTOCs. Specifically, by using classical tensor network methods to generate a small amount of training data, training a model using the RBF or Laplacian kernel, then applying the model to make predictions, we can quickly generate accurate approximations of OTOCs. One must, of course, incur the cost of producing the training data, which can be expensive. However, in cases where one needs to evaluate OTOCs for many different systems and values of time belonging to the short- and possibly intermediate-time regimes, we can use kernel methods to reduce the cost of many of the predictions to a time scaling linearly in the number of training datapoints, as demonstrated by (36).
@@
The paper naturally lends itself to a handful of possible extensions, which we suggest as future research directions. For example, one may like to investigate whether the KM approach works with other OTOCs, other linear combinations of OTOCs, or other types of correlation functions such as the Loschmidt echo [83] or entanglement entropy [60]. We could also examine how the approach works for other parameterised sets of Hamiltonians. For instance, we could investigate sets with $d>3$ (i.e., parameterised by four or more parameters), sets containing Hamiltonians describing quantum systems in higher dimensions, or sets containing highly non-local Pauli strings in one or more terms. Though producing data for the latter two cases would likely be challenging. It might also be interesting to apply quantum kernel methods [84, 85, 86] to investigate whether a problem-specific quantum kernel may offer significant increases in learning performance. Similarly, it would be useful to investigate how the results change in response to increasing the size of the parameter region from which we sample the input data. This may shed light on whether the kernels can be used to accurately predict long-time OTOCs, though producing the data in this case would likely be expensive. Finally, as discussed earlier, it would be useful to explore whether a predictable trend in the hyperparameter values may be uncovered by using a more finely grained grid search.
@@
The datasets produced in this study are available in the GitHub repository at https://github.com/John-J-Tanner/ITensor-OTOC-Data. The repository includes all relevant datasets necessary to replicate the study’s results, and example Python codes that can be used to visualise the data. For further information or inquiries, please contact the corresponding author, John Tanner, at john.tanner@uwa.edu.au.
@@ Appendix A Supplementary results
Appendix B Convergence in the bond dimension
In this appendix, Figure B.1 provides insight into how well the labels in the datasets produced in this work have converged with respect to the bond dimensions used to calculate them. These results are discussed in more detail in the second last paragraph of Sec. V.
@@ Appendix B Convergence in the bond dimension
Figure B.1: The numerical values of $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ for ten randomly sampled inputs plotted against the maximum bond dimension $\chi$ used to calculate them for 40-qubit quantum systems. Plots (a), (b), (c) and (d) show $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Plots (e), (f), (g) and (h) show $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. The vertical black lines indicate the value of $\chi$ used to calculate the rest of the data for the corresponding function and parameterised set of Hamiltonians. In the same order as the plots, the chosen values of $\chi$ are 110, 140, 110, 150, 70, 90, 70 and 100.
Appendix C Hyperparameter values
In this section of the appendices, Table C.1 provides a list of the hyperparameter values which were trialled for each kernel during the 10-fold cross-validation. Of all the values which were trialled, Tables C.2, C.3, C.4, and C.2 then provide the best hyperparameter values which achieved the highest average coefficient of determination over all ten folds of the training dataset. These hyperparameter values were then used to train the kernels on the full training datasets and subsequently apply the trained models to the testing data.
[…]
Table C.1: Lists of hyperparameters values which were trialled during the 10-fold cross-validation performed on the training datasets. The hyperparameter $\lambda$ refers to the regularisation strength which appears in (1), while the hyperparameters $\gamma$, $c_{0}$ and $d$ correspond to those which appear in definitions of the kernels in (26), (27), (28), (29), (30) and (31).
@@ Appendix C Hyperparameter values
Table C.2: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{1}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all ten folds of the training dataset.
@@
Table C.3: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{2}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all ten folds of the training dataset.
@@
Table C.4: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{3}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all ten folds of the training dataset.
@@
Table C.5: The best hyperparameters trialled during the 10-fold cross-validation performed on the training datasets for $\mathscr{H}_{4}$. Of all the hyperparameters trialled, the ones listed here resulted in models which achieved the highest average coefficient of determination ($R^{2}$) over all ten folds of the training dataset.
Appendix D Numerical results
The tables contained in this section of the appendices contain the majority of the numerical results reported in this work. Here we provide learning performance metrics for the models trained on all 1000 training datapoints with the hyperparameter values listed in the last four tables of Appendix C. Specifically, for every choice of $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$ with one of the parameterised sets of Hamiltonians $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$, or $\mathscr{H}_{4}$, we provide the performance metrics calculated based on predictions made by the trained models on both the training and testing sets.
D.1 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{1}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$1.9284 -4.1339 -3.5942 -3.7616 -3.4297 -3.3197 -3.0432 -2.8091 RMSE 0.7886 0.8122 0.8156 0.8214 0.8156 0.8147 0.8030 0.7941 MAE 0.7091 0.7296 0.7275 0.7307 0.7186 0.7159 0.6991 0.6857 Polynomial $R^{2}$ 0.8604 0.8265 0.7872 0.8276 0.8145 0.8132 0.8154 0.8097 RMSE 0.1722 0.1493 0.1755 0.1563 0.1669 0.1694 0.1716 0.1775 MAE 0.1228 0.1076 0.1185 0.1121 0.1200 0.1191 0.1249 0.1274 RBF $R^{2}$ 0.9646 0.9454 0.8910 0.8963 0.8771 0.9038 0.9111 0.9246 RMSE 0.0867 0.0837 0.1256 0.1212 0.1359 0.1216 0.1191 0.1117 MAE 0.0431 0.0489 0.0586 0.0772 0.0790 0.0739 0.0674 0.0594 Laplacian $R^{2}$ 0.9391 0.9696 0.9228 0.9547 0.9356 0.9362 0.9401 0.9542 RMSE 0.1137 0.0625 0.1057 0.0801 0.0983 0.0990 0.0978 0.0871 MAE 0.0763 0.0388 0.0495 0.0463 0.0529 0.0660 0.0636 0.0559 Sigmoid $R^{2}$ 0.6777 0.8639 0.6475 0.8191 0.8144 0.8461 0.6852 0.6400 RMSE 0.2616 0.1322 0.2259 0.1601 0.1669 0.1538 0.2240 0.2441 MAE 0.1852 0.0895 0.1617 0.1123 0.1176 0.1081 0.1696 0.1899 Cosine $R^{2}$1.9508 -4.1287 -3.5710 -3.7659 -3.4271 -3.3113 -3.0796 -2.8382 RMSE 0.7916 0.8117 0.8135 0.8218 0.8153 0.8139 0.8066 0.7971 MAE 0.7116 0.7285 0.7240 0.7306 0.7175 0.7136 0.7024 0.6882 Training Linear $R^{2}$1.7147 -3.4479 -3.1980 -3.1767 -2.9977 -2.8207 -2.6211 -2.4198 RMSE 0.7692 0.7919 0.7971 0.8052 0.8025 0.7979 0.7870 0.7745 MAE 0.6823 0.6985 0.6996 0.7035 0.6971 0.6893 0.6750 0.6589 Polynomial $R^{2}$ 0.8942 0.8747 0.8709 0.8712 0.8715 0.8779 0.8650 0.8572 RMSE 0.1519 0.1329 0.1398 0.1414 0.1439 0.1426 0.1520 0.1583 MAE 0.1071 0.0939 0.0992 0.1006 0.1035 0.0997 0.1078 0.1110 RBF $R^{2}$ 0.9997 0.9921 0.9987 0.9578 0.9724 0.9686 0.9803 0.9871 RMSE 0.0077 0.0334 0.0138 0.0809 0.0667 0.0723 0.0580 0.0476 MAE 0.0049 0.0224 0.0078 0.0557 0.0452 0.0491 0.0390 0.0308 Laplacian $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9998 0.9999 RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0075 0.0059 0.0047 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0054 0.0042 0.0033 Sigmoid $R^{2}$ 0.7042 0.9410 0.7206 0.8773 0.8743 0.9130 0.7229 0.6310 RMSE 0.2539 0.0912 0.2056 0.1380 0.1423 0.1204 0.2177 0.2544 MAE 0.1791 0.0666 0.1470 0.0996 0.1009 0.0881 0.1567 0.1892 Cosine $R^{2}$1.7129 -3.4530 -3.2072 -3.1761 -2.$9$2.8250 -2.6125 -2.4122 RMSE 0.7690 0.7923 0.7979 0.8052 0.8027 0.7983 0.7860 0.7736 MAE 0.6816 0.6983 0.6983 0.7030 0.6961 0.6873 0.6746 0.6578
@@
Table D.1: ($\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.2. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.2 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{2}$